OpenStudy (anonymous):

If a^2-5a+1=0 then find the value of 1) a^3+a^-3 2)a^2+a^-2

OpenStudy (anonymous):

Find the value of 'a' from the first equation by factorising or quadratic formula than sub... it in the next two equations

OpenStudy (anonymous):

well that is one way. I'm sure there is a better way to solve the problem, probably using some algebraic formulas.

OpenStudy (loser66):

I go nowhere, hihihi,,, but just want to give out my idea \(a^2 = 5a -1\\a^3 = 5a^2 -a\) and combine them as you wish.

OpenStudy (anonymous):

where do you go after that?

OpenStudy (loser66):

Is it not that \(a^3+a^{-3} = a^3 + \dfrac{1}{a^3}\)??

OpenStudy (anonymous):

exactly it is.

OpenStudy (loser66):

so, just put my a^3 in, hihihi

OpenStudy (anonymous):

i understood that. The problem arises when I try to simplify things beyond that. Any clues on that?

ganeshie8 (ganeshie8):

maybe we can try using symmetry to ease arithmetic : just notice that if \(\large \alpha \) is a root of given quadratice, then \(\large \dfrac{1}{\alpha }\) is also a root.

OpenStudy (anonymous):

ok.. go on

OpenStudy (anonymous):

or use a^3 + b^3

ganeshie8 (ganeshie8):

next observe that \(\large a^2 + a^{-2} = \left(a + a^{-1} \right)^2 - 2 \)

OpenStudy (anonymous):

(a+b)^3-3ab(a+b)

OpenStudy (anonymous):

yes that is quite right.

OpenStudy (anonymous):

you know sum and product of roots , by vieta's formula

OpenStudy (anonymous):

that is , -b/a and c/a

OpenStudy (anonymous):

well , i have not tried , maybe 90% it would work

ganeshie8 (ganeshie8):

brilliant ! you have just read out my mind :) it works nicely

OpenStudy (anonymous):

Eureka ! thanks

ganeshie8 (ganeshie8):

say the roots of given quadratic are \(\large a, a^{-1}\) : \(\large a + a^{-1} = 5\) \(\large aa^{-1 } = 1\) \(\large \implies a^2 + a^{-2} = \left(a + a^{-1} \right)^2 - 2 = 5-1 = 4 \) you can work out \(\large a^3+a^{-3}\) as suggested by @No.name

OpenStudy (anonymous):

Well if you know vieta's relation(this relation is different) , you can very well compute a^10 also i learnt the method a couple of does back it is really a nice one

ganeshie8 (ganeshie8):

*\(\large \implies a^2 + a^{-2} = \left(a + a^{-1} \right)^2 - 2 = 5-2 = 3\)

OpenStudy (anonymous):

It is similar to theory of equations well @deeprony7 you got it

OpenStudy (anonymous):

would you mind sharing the vieta's relation that you mentioned? @No.name

OpenStudy (anonymous):

i watched a video , i will share , give me a moment

OpenStudy (anonymous):

the ans to a^2+a^-2 is actually 23(atleast thats what it says in my book) @ganeshie8

OpenStudy (anonymous):

https://www.youtube.com/watch?v=y8sQBGK9yYM

OpenStudy (anonymous):

Well that is a cubic

ganeshie8 (ganeshie8):

Ahh yes, was just checking whether you're really paying attention or not :P there is a mistake in calculation, haven't u noticed i forgot to square

ganeshie8 (ganeshie8):

corrected below\[\large \implies a^2 + a^{-2} = \left(a + a^{-1} \right)^2 - 2 = 5^2-2 =2 3\]

OpenStudy (anonymous):

You may not learn that method now

ganeshie8 (ganeshie8):

I think that video is talking about Newton's symmetric sums http://www.artofproblemsolving.com/Wiki/index.php/Newton's_Sums

OpenStudy (anonymous):

Oh , someone told me its some vieta thing , lol

OpenStudy (anonymous):

It's Newton's sum i c